Circular No. 7457

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

NOVA IN THE LARGE MAGELLANIC CLOUD 2000 H. W. Duerbeck, Vrije Universiteit, Brussels; and E. Pompei, European

Southern Observatory (ESO), write: “A spectrum (range 385–680 nm) of this nova (IAUC 7453) that was taken with the 1.5-m Danish telescope (+ DFOSC) at ESO, La Silla, on July 15.4 UT shows emission lines of the Balmer series, Fe II multiplets 42, 37, 38, and 74, and Na I D, superimposed on a weak continuum, indicating a ‘Fe II’ nova about a week after maximum. The Balmer lines have FWHM = 1700 km/s and very weak absorption components at −1900 km/s relative to the line centers.”

J. B. Hearnshaw and J. Yan Tse, University of Canterbury, report medium-resolution CCD spectroscopy (range 380–680 and 800–910 nm) of this nova on July 14 and 15 UT using the 1-m telescope at Mount John University Observatory. Balmer lines Hα to H_{9} were all prominent. Other emission lines were Fe I (490.3, 500.6, and 541.5 nm), Mg I (518.4 nm), Fe II (531.7 nm), Na D (589.3 nm), O I (844.6 nm), and several Paschen lines (P_{9}, P_{10}, P_{11}). The FWZI of the Balmer lines is ∼ 2200 ± 250 km/s. P-Cyg absorption components were not found. I. A. Bond and P. M. Kilmartin find that the nova was not present on CCD images taken with the Mount John 0.6-m telescope on June 29.35, but it was clearly visible on June 29.68. The Mount John image showing the nova at its brightest was obtained on July 1.68. Magnitude calibrations are being made.

Additional broadband V magnitudes from W. Liller, Viña del Mar, Chile: July 13.9762 UT, 11.40; 15.9645, 11.72.

SUPERNOVA 2000cw IN MCG +5-56-007 M. Dennefeld, Institut d’Astrophysique de Paris, reports on behalf a

collaboration including R. Gredel, Heidelberg; A. Pizzella, Padova; P. Leisy, European Southern Observatory; A. Budovičová, Prague; M. Markoulaki, Athens; I. Pascucci, Jena; and himself: “We obtained a low-dispersion spectrum (resolution 0.45 nm, range 490–820 nm) of SN 2000cw (cf. IAUC 7456) on July 17.161 UT with the 2.2-m telescope (+ CAFOS) at Calar Alto Observatory. The spectrum shows S II at 545.0–560.0 nm, Si II at 597.0 nm, and the strong Si II 635.5-nm absorption dips (measured at 544, 560, 594, and 629 nm, respectively) that are characteristic of a type-Ia supernova within a few days of maximum. From the 635.5-nm feature, with a heliocentric velocity of 9030 km/s for the galaxy (Lyon database), we deduce an expansion velocity of 11 300 km/s.”

2000 July 17  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Daniel W. E. Green